# Diffusing DeBias: Synthetic Bias Amplification for Model Debiasing

### v3

@@ 4.4 Ablation analysis
In this section, we provide an extensive ablation analysis of DDB main components. Without losing generality, all ablations are performed on Waterbirds and exploit Recipe I for model debiasing. Additional ablations on Recipe II can be found in the Suppl. Material. The reported analyses regard several aspects of the data generation process with respect to the Bias Amplifier, the effects of the filtering option in Recipe I, and the behavior of DDB on unbiased data.
Bias amplifier data requirements. All our main results are obtained by training a BA on $1000$ synthetic images per class. Here, we want to measure how this choice impacts the effectiveness of our Bias Amplifier. From Table 2 (left), we observe that even 100 synthetic images are enough for good debiasing performance despite not being competitive with top accuracy scores. As we increase the number of training images, WGA improves, reaching the best performance for $2000$ images per class.
CFG strength and generation bias. The most important free parameter for the CDPM with CFG is the guidance conditioning strength $w$. Such strength can affect the variety and consistency of generated images. Here, we investigate DDB dependency on this parameter. Table 2 (right) shows a limited impact on debiasing performance. WGA variations along different $w$ values are limited, with the best performance obtained for $w=1$.
@@
Group extraction via BA plain error sets. In Sec. 3.3.1, we describe how we filter the BA error set through a simple heuristic based on per-sample loss distribution. In Table 1, the last row (DDB-I w/ plain err. set) reports our obtained results when considering as bias-conflicting all the samples misclassified by the Bias Amplifier. Our results show a trade-off between the two alternatives. The entire set of misclassifications are always enough to reach high bias mitigation performance with less critical settings ($\rho=0.950$) (e.g., Waterbirds), gaining an advantage from such a coarser selection. On the other hand, the importance of a more precise selection is highlighted by the superior performance of the filtered alternative in the more challenging scenarios, like BFFHQ ($\rho=0.995$).
Bias-Identification Accuracy To frame the goodness of our Bias Amplifier in correctly identifying aligned and conflicting training samples, we provide the accuracy of such process when considering samples misclassified by the BA. Specifically, we report a bias-identification accuracy of $89.00\%$ for Waterbirds and $96.00\%$ on BFFHQ. Our very high accuracy supports the effectiveness of using the proposed protocol, which is not dependent from any careful regularization, bias-annotated validation sets, or large ensembles of auxiliary classifiers.
DDB impact on unbiased dataset. In real-world settings, it is generally unknown whether bias is present in datasets or not. Consequently, an effective approach, while mitigating bias dependency in the case of a biased dataset should also not degrade the performance when the bias is not present. This ablation study aims to assess DDB from this perspective. To obtain this insight, we apply the entire DDB pipeline (adopting Recipe I) on a common image dataset such as CIFAR-10  [46], comparing it with traditional ERM training with CE loss. In this controlled case, Recipe I provides a test accuracy of $84.16\%$, with a slight improvement of $\sim 1\%$ compared to traditional ERM ($83.26\%$) employing the same target model (ResNet18). This result shows that our Bias Amplifier can still provide beneficial signals to the target model in standard learning scenarios, potentially favoring the learning of the most challenging samples.
@@ 5 Conclusions
In this work, we present Diffusing DeBias (DDB), a novel debiasing framework capable of learning the per-class bias-aligned training data distribution, leveraging a CDPM. Synthetic images sampled from the inferred biased distribution are used to train a Bias Amplifier, employed to provide a supervisory signal for training a debiased target model. The usage of synthetic bias-aligned data permits avoidance of detrimental effects typically impacting auxiliary models’ training in unsupervised debiasing schemes, such as bias-conflicting sample overfitting and interference, which lead to suboptimal generalization performance of the debiased target model. DDB is effective and versatile, acting as a plug-in for unsupervised debiasing methods. In this work, we design a two-step and an end-to-end debiasing recipes incorporating DDB’s Bias Amplifier: they outperform state-of-the-art works by a margin on four popular biased benchmark datasets, including the most recent relying on vision-language models [23, 47]. Finally, while effectively mitigating bias dependency for biased datasets, DDB does not degrade performance when used with unbiased data, making it suitable for debiasing in real-world applications.
@@
Appendix
Appendix A Ablation Studies with DDB’s Recipe II
@@
Appendix B Full Implementation Details of Recipe I (two-step)
Appendix C Full Implementation Details of Recipe II (end-to-end)
For BAR, we use the same parameters reported in [6]: batch size = 128, learning rate = $1\times e^{-4}$, weight decay = 0.01, training epochs = 50. Waterbirds: batch size = 128, learning rate = $5\times e^{-5}$, weight decay = 0.01, training epochs = 80. BFFHQ: batch size = 256, learning rate = $5\times e^{-5}$, weight decay = 0.01, training for 50 epochs. For these recipes, we accumulate gradients for $8$ iterations ($16$ for BFFHQ) with mini-batches made of $16$ samples.
Appendix D Additional Biased Generations
@@
Appendix E Bias Validation via Captions in Synthetic Images
As an ulterior confirmation of bias in CDPM-generated images, besides two independent annotators, we implement an unsupervised analysis pipeline using the pre-trained BLIP-2 FlanT5XL model [40] for zero-shot image captioning without any task-specific prompting, for avoiding any guidance in the descriptions of the synthetic data. Our method processes the resulting captions through stop word removal and frequency analysis to reveal underlying biases without relying on predetermined categories or supervised classification models. Figure 10 presents the keyword frequency distribution for the Waterbirds dataset, where the most prevalent terms naturally correspond to known bias attributes (e.g., environmental contexts). The same is true for the keyword frequency distribution of the BFFHQ dataset (Figure 11) and BAR (Figure 12) where the most prevalent terms are related to the gender and the context respectively. These findings confirm that generated images are bias-aligned. Furthermore, the obtained word frequency histograms suggest that images generated by DDB’s bias diffusion step may provide information useful to support bias identification (or discovery) at the dataset level.
@@ Appendix E Bias Validation via Captions in Synthetic Images
Figure 11: Word frequency analysis from 1000 generated captions for BFFHQ classes ‘young’ (left) and ‘old’ (right), showing top 10 most frequent terms.
(a) Class 0
(b) Class 1
(c) Class 2
(d) Class 3
(e) Class 4
(f) Class 5
Figure 12: Word frequency analysis from 1000 generated captions for BAR classes, showing top 10 most frequent terms for each class.

### v4

@@ 4.3 Ablation analysis
In this section, we provide an extensive ablation analysis of DDB main components. Without losing generality, all ablations are performed on Waterbirds and exploit Recipe I for model debiasing. Additional ablations on Recipe II can be found in the Supp. Mat. The reported analyses regard several aspects of the data generation process with respect to the Bias Amplifier, the effects of the filtering option in Recipe I, and the behavior of DDB on unbiased data.
Bias Amplifier data requirements. All our main results are obtained by training a BA on $1000$ synthetic images per class, and we want here to measure how this choice impacts its effectiveness. From Table 3 (left), we observe that even 100 synthetic images are enough for good debiasing performance despite not being competitive with top accuracy scores. Increasing the number of training images improves WGA, reaching the best performance for $1000$ images per class.
CFG strength and generation bias. The most important free parameter for the CDPM with CFG is the guidance conditioning strength $w$. Such strength can affect the variety and consistency of generated images. Here, we investigate DDB dependency on this parameter. Table 3 (right) shows a limited impact on debiasing performance. WGA variations along different $w$ values equal or larger than 1 are limited, with the best performance obtained for $w=1$.
@@
Group extraction via BA plain error sets. In Sec. 3.3.1, we describe how we filter the BA error set through a simple heuristic based on per-sample loss distribution. In Table 1, the last row (DDB-I w/ plain err. set) reports our obtained results when considering as bias-conflicting all the samples misclassified by the Bias Amplifier. Our results show a trade-off between the two alternatives. The entire set of misclassifications are always enough to reach high bias mitigation performance with less critical settings (e.g., like in Waterbirds, $\rho=0.950$), gaining an advantage from such a coarser selection. On the other hand, the importance of a more precise selection is highlighted by the superior performance of the filtered alternative in the more challenging scenarios, like in BFFHQ ($\rho=0.995$).
Bias-Identification Accuracy. To frame the goodness of our BA in correctly identifying aligned and conflicting training samples, we provide the accuracy of such a process when considering samples misclassified by it. Specifically, we reach a high bias-identification accuracy of $89.00\%$ for Waterbirds and $96.00\%$ on BFFHQ, supporting the effectiveness of the proposed protocol, which is not dependent on any careful regularization, bias-annotated validation sets, or large ensembles of auxiliary classifiers.
DDB impact on unbiased dataset. In real-world settings, it is generally unknown whether bias is present in datasets or not. Consequently, an effective approach, while mitigating bias dependency in the case of a biased dataset, should also not degrade the performance when the bias is not present. To evaluate DDB in this regard, we apply the entire pipeline (adopting Recipe I) on a common image dataset such as CIFAR-10  krizhevsky2009learning , comparing it with traditional ERM training with CE loss. In this controlled case, Recipe I provides a test accuracy of $84.16\%$, with a slight improvement of $\sim 1\%$ compared to traditional ERM ($83.26\%$) employing the same target model (ResNet18). This result shows that our Bias Amplifier can still provide beneficial signals to the target model in standard scenarios, potentially favoring the learning of the most complex samples.
@@ 5 Conclusions
In this work, we present Diffusing DeBias (DDB), a novel debiasing framework capable of learning the per-class bias-aligned training data distribution, leveraging a CDPM. Synthetic images sampled from the inferred biased distribution are used to train a Bias Amplifier, employed to provide a supervisory signal for training a debiased model. The usage of synthetic bias-aligned data prevents the detrimental effects typically impacting auxiliary models’ training in unsupervised debiasing schemes, such as bias-conflicting sample overfitting and interference, which lead to suboptimal generalization performance of the debiased target model. DDB results effective and versatile, acting as a plug-in for unsupervised debiasing methods. In this work, we incorporate DDB’s Bias Amplifier in a two-step and an end-to-end debiasing recipes. leading to accuracies outperforming state-of-the-art works by a margin on popular biased benchmark datasets (considering single and multiple biases), including recent works relying on vision-language models kim2024discovering ; ciranni2024say . Finally, while effectively mitigating bias dependency for biased datasets, DDB does not degrade performance when used with unbiased data, making it suitable to be adopted as an effective general-purpose standard training procedure.
@@
Appendix A Technical Appendices and Supplementary Material
A.1 Datasets
  • Waterbirds  Sagawa*2020Distributionally is an image dataset exhibiting strong correlations ($\rho=0.950$) between bird species and background environments (e.g., water vs. land). It has 4,795 training images, 1,199 images for validation, and 5,794 for testing.
[…]
  • We utilize the original BAR version as introduced in nam2020learning . The dataset presents 6 different classes, but does not provide explicit bias annotations or a validation set, with 1,941 images for training and 654 for testing.
  • ImageNet-A consists of 7,500 real-world images from a 200-class subset of ImageNet, where deep learning models consistently struggle to make correct predictions. Originally introduced by Hendrycks et al. hendrycks2021natural to evaluate adversarial robustness, it is also employed as a model debiasing benchmark, usually as a bias-conflicting evaluation set for models trained on ImageNet-9 bahng2019rebias , a collection of images with 57,200 samples where categories are super-classes of ImageNet, put together to exhibit textural and shape biases bahng2019rebias . As such, both our CDPM and Bias Amplifier are trained on ImageNet-9. Then, we evaluate both recipes on ImageNet-A in terms of Average Accuracy, following the same protocol of bahng2019rebias ; NEURIPS2022_75004615_LWBC ; nahon2023mining .
  • UrbanCars has been released in li2023whac and contains 8000 training images with the target represented by urban and country cars. The dataset is built starting from the original $196$ classes of Stanford Cars stanfordcars , while synthetically incorporating two different shortcuts, background, from the Places dataset zhou2017places , and Co-Occurring Objects (CoObj), from the LVIS one gupta2019lvis . In UrbanCars, $90.25\%$ of samples present both background and CoObj attributes aligned with the target class biases, $9.5\%$ of training samples are conflicting with respect to only one of the two attributes, while $0.25\%$ are conflicting to both. The evaluation metrics employed for this datasets are the one introduced in li2023whac , namely ID.ACC, BG-Gap, CoObj-Gap, BGCoObj-Gap. ID.ACC is the in-distribution accuracy (the weighted per-class accuracy using the training set per-group subpopulation proportions). The other metrics represent the test accuracy gap of each minority subgroup with respect to ID.ACC.
A.2 Implementation Details
In our experiments, we generate $1000$ synthetic images for each target class using a CDPM implemented as a multi-scale UNet incorporating temporal and conditional embeddings for processing timestep information and class labels. For computational efficiency, input images are resized to $64\times 64$ pixels. Input images are normalized with respect to ImageNet’s mean and standard deviation. The CDPM is trained from scratch for $70,000$, $150,000$, $200,000$, $100.000$, $100.000$ iterations for the datasets BAR, Waterbirds, BFFHQ, ImageNet-9, and UrbanCars respectively. Batch size is set to $32$, and the diffusion process is executed over $1,000$ steps with a linear noise schedule (where $\beta_{1}=1e^{-4}$ and $\beta_{T}=0.028$). Optimization is performed using MSE loss and AdamW optimizer, with an initial learning rate of $1e^{-4}$, adjusted by a CosineAnnealingLR DBLP:conf/iclr/LoshchilovH17 scheduler with a warm-up period for the first $10\%$ of the total iterations. As per the Bias Amplifier training, we use synthetic images from CDPM. A Densenet-121 huang2017densely with ImageNet pre-training is employed across all datasets, featuring a single-layer linear classification head. The training uses regular Cross-Entropy loss function and the AdamW optimizer loshchilov2018decoupled . For BFFHQ, BAR, and UrbanCars, the learning rate is $1e^{-4}$, and $5e^{-4}$ for Waterbirds, and ImageNet-9. Training spans 50 epochs with weight decay $\lambda=0.01$ except for BFFHQ, which uses $\lambda=1.0$ for 100 epochs. We apply standard basic data augmentation strategies (following kim2021biaswap ; nam2020learning ), including random crops and horizontal flips. For our debiasing recipes, we rely on the same backbones used in the existing literature, i.e.ResNet-18 for BFFHQ, BAR, ImageNet-9/A, and a ResNet-50 for Waterbirds and UrbanCars nam2020learning ; kim2021biaswap ; kim2024discovering ; li2023whac . In Recipe II, $\gamma$ (Equation 3) is set to 3 in all our main experiments. Finally, similarly to li2022discover ; nam2020learning ; pastore2024lookingmodeldebiasinglens , we do not rely on a validation set for regularization. This choice is related to real-world application suitability, where no guarantees exist regarding the degree of spurious correlations present in a held-out subset of data used as validation, possibly resulting in detrimental regularization zarlenga2024efficient .
A.3 Full Implementation Details of Recipe I (two-step)
For UrbanCars, presenting two biases, we apply Recipe I as it is, considering the filtered correct and wrong classifications in each class as 4 subgroups during GDRO optimization. We then consider all the available 8 subgroups at evaluation time.
A.4 Full Implementation Details of Recipe II (end-to-end)
For BAR, we use the same parameters reported in nam2020learning : batch size = 128, learning rate = $1\times e^{-4}$, weight decay = 0.01, training epochs = 50. Waterbirds: batch size = 128, learning rate = $5\times e^{-5}$, weight decay = 0.01, training epochs = 80. BFFHQ: batch size = 256, learning rate = $5\times e^{-5}$, weight decay = 0.01, training for 50 epochs. For these recipes, we accumulate gradients for $8$ iterations ($16$ for BFFHQ) with mini-batches made of $16$ samples.
A.5 Ablation Studies with DDB’s Recipe II
@@
A.6 Additional Biased Generations
@@
Figure 5: Synthetic bias-aligned image generations for each class of the UrbanCars dataset. Each grid displays 100 synthetic images per class, highlighting model bias alignment.
@@
A.7 Bias Validation via Captions in Synthetic Images
As an ulterior confirmation of bias in CDPM-generated images, besides two independent annotators, we implement an unsupervised analysis pipeline using the pre-trained BLIP-2 FlanT5XL model DBLP:conf/icml/0008LSH23 for zero-shot image captioning without any task-specific prompting, for avoiding any guidance in the descriptions of the synthetic data. Our method processes the resulting captions through stop word removal and frequency analysis to reveal underlying biases without relying on predetermined categories or supervised classification models. Figure 9 presents the keyword frequency distribution for the Waterbirds dataset, where the most prevalent terms naturally correspond to known bias attributes (e.g., environmental contexts). These findings confirm that generated images are bias-aligned. Furthermore, the obtained word frequency histograms suggest that images generated by DDB’s bias diffusion step may provide information useful to support bias identification (or discovery) at the dataset level.